y= -x/2-4
-2x-y=-5

Answer by jim_thompson5910(35256)

Hint:

y= -x/2-4

2y = -x - 8

-x = 2y + 8

x = -2y - 8

-------------------------------------------------------

-2x-y=-5

-2(-2y - 8)-y=-5

4y + 16 - y = -5

Solve this equation for y. Then use it to find x.